1. What is the molecular mass of a non-electrolyte solute in a solution that contains 0.6kg of sol 200g of benzene with a freezing point equal to 5.4°C? The freezing point of pure benzene is 5.5 The freezing point constant of benzene is 5.12°C.

1
Expert's answer
2022-03-25T07:54:04-0400

∆T = Kf × m

∆T = Tpure - Tsol = 5.5 - 5.4 = 0.1°C

0.1°C = 5.12°C/m × m

m = 0.1/5.12 = 0.0195 mol/kg

m = Moles of solute / kg of solvent

m = m(solute) /(Mr(solute) × kg (solvent)

0.0195 mol/kg = 200 g /(Mr(solute) × 0.6 kg)

Mr(solute) = 200 g /(0.0195 mol/kg × 0.6 kg) = 17 094 g/mol